Determine the seating capacity of an auditorium with 18 rows of seats if there are 20 seats in the first row, 24 seats in the second row, and 28 seats in the third row and so on
Answer by richwmiller(17219)


S = ½(2a + (n-1)d)n
S = ½(40 + 4*17)18
S = 9(40 + 68)
s=9*108=972 seats in 18 rows
row seats total
1 20 20
2 24 44
3 28 72
4 32 104
5 36 140
6 40 180
7 44 224
8 48 272
9 52 324
10 56 380
11 60 440
12 64 504
13 68 572
14 72 644
15 76 720
16 80 800
17 84 884
18 88 972